OpenStudy (anonymous):

Simplify the expression 63 + 5(4 − 2). 28 36 226 234

OpenStudy (anonymous):

Medal and fan :) @mathway @Zale101 @Mertsj @Lady.Liv1776 @YanaSidlinskiy @PhoenixBirrd

OpenStudy (mertsj):

What is 4-2?

OpenStudy (anonymous):

2

OpenStudy (mertsj):

Replace 4-2 with 2

OpenStudy (anonymous):

Ok

OpenStudy (mertsj):

What is 5(2)?

OpenStudy (anonymous):

10

OpenStudy (lynfran):

r u sure u wrote the expression correct?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

@mathway

OpenStudy (anonymous):

@pooja195

pooja195 (pooja195):

Use order of operations

pooja195 (pooja195):

PEMDAS Parentheses Exponents Multiplication Division Addition Subtraction

pooja195 (pooja195):

Basically what @Mertsj was telling you to do : )

pooja195 (pooja195):

And something is wrong....

OpenStudy (lynfran):

yep

OpenStudy (anonymous):

Really?

OpenStudy (lynfran):

yes 63+10=73 and...73 are not in ur answer options...

OpenStudy (anonymous):

Oh yeah @pooja195 @LynFran 3 is an exponent sorry i didnt notice

OpenStudy (lynfran):

ok ...well what 6*6*6=?

OpenStudy (anonymous):

216

pooja195 (pooja195):

\[\huge~\rm~ 6^3 + 5(4 − 2).\]

OpenStudy (lynfran):

good so 216+10=?

OpenStudy (anonymous):

226 Thank you :)

OpenStudy (lynfran):

welcome